Two objects of equal masses collide head-on while moving in the opposite direction and joined together. Show that the fractional loss of original kinetic energy of the objects is given by the expression; E=1/2+V1V2/V1²+V2²

1
Expert's answer
2022-03-07T11:45:17-0500

Explanations & Calculations


  • Final speed needs to be found using the conservation of linear momentum.

"\qquad\qquad
\begin{aligned}
\small \to\\
\small mv_1-mv_2&=\small 2mV\\
\small V&=\small \frac{v_1-v_2}{2}
\end{aligned}"

  • Then the energy loss is

"\qquad\qquad
\begin{aligned}
\small L&=\small K.E_{inital}\\
\small \Delta L&=\small K.E_{inital}-K.E_{final}\\
&=\small \frac{1}{2}mv_1^2+\frac{1}{2}mv_2-\frac{1}{2}.2m.V^2\\
&=\small \frac{m}{2}\Big[\frac{v_1^2}{2}+\frac{v_2^2}{2}+v_1v_2\Big]\\

\end{aligned}"

  • Then the fractional loss is

"\qquad\qquad
\begin{aligned}
\small \frac{\Delta L}{L}&=\small \frac{\frac{v_1^2}{2}+\frac{v_2^2}{2}+v_1v_2}{v_1^2+v_2^2}\\
&=\small \frac{\frac{(v_1^2+v_2^2)}{2}+v_1v_2}{(v_1^2+v_2^2)}\\
&=\small \frac{1}{2}+\frac{v_1v_2}{(v_1^2+v_2^2)}
\end{aligned}"